# The Sociotechnical Stack: Opportunities for Social Computing Research in Non-consensual Intimate Media

### v1

@@ Abstract.
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2024; Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Collaborative and social computing theory, concepts and paradigms
@@ 1. Introduction
Non-consensual intimate media (NCIM) is the unauthorized creation, sharing, or distribution of sexual content containing someone’s body or likeness. This form of online abuse usually targets women, who constitute about 90% of all cases (Eaton et al., 2017; Initiative, 2014; saf, [n. d.]). NCIM cases can be varied, ranging from sexually explicit deepfakes, to covert recordings, to sextortion (See Fig. 2). NCIM is common—studies suggest 1 in 6 adults have had their intimate content shared without consent (Powell et al., 2024; Ruvalcaba and Eaton, 2020). NCIM is also deeply traumatic—almost all victim-survivors experience severe anxiety and depression. About half seriously consider ending their own lives, and in some cases women and girls commit suicide to escape emotional pain and social shame. (Compton and Hunt, 2024; Initiative, 2014; Eaton et al., 2017). While there is widespread agreement that NCIM is harmful and some agreement that it should be regulated, there is less consensus on how to do so (Citron and Franks, 2014; De Angeli et al., 2023; Citron and Franks, 2020; Franks, 2014). The law is evolving to recognize NCIM, but access to legal recourse is limited and cannot prevent continued re-distribution. Victim-survivors may request content removal directly from platforms, but these efforts are typically laborious, emotionally taxing, and frequently futile due to the diverse jurisdictions governing online platforms (McGlynn et al., 2021; De Angeli et al., 2023).
@@
This paper highlights the essential links between NCIM and its enabling technical components. Drawing on existing surveys and interviews with more than 400 victim-survivors, we identify a critical gap: technologies, including social media platforms, file formats, and content recommendation algorithms, do not merely enable but actively create NCIM at scale. To systemically analyze this, we introduce a conceptual framework called the sociotechnical stack. In software engineering and computer science, a “tech stack” refers to a layered structure of technologies used to build systems, like the TCP/IP protocols in computing networking. We extend this concept to include social impacts, integrating the technical and social dimensions of systems. Our analysis shows how certain layers, such as a content recommendation algorithm, can unintentionally promote NCIM. The sociotechnical stack illustrates the ways our digital environment facilitates risks and harms for NCIM victim-survivors, including the non-consensual creation of sexual content via generative AI and unauthorized redistribution of privately shared intimate media. We propose a roadmap for sociotechnical research to address NCIM alleviate risks for victim-survivors.
@@
  • The sociotechnical stack conceptual framework: Intended for analyzing social computing issues, the sociotechnical stack framework breaks down complex sociotechnical problems onto specific technical components that contribute to social impacts and harms.
@@ 2. Overview of NCIM
NCIM’s prevalence is extensive and likely underreported, compounded by technical challenges in detection and social stigma that deter reporting. One 2017 survey found that 5% of respondents admitted to non-consensually sharing intimate photos (Eaton et al., 2017). This almost certainly under-counts people who consume NCIM online. A 2014 study estimated around 11,000 websites hosted non-consensual content—a figure that has likely grown (Citron, 2014; Initiative, 2014; Eaton et al., 2017). These sites range in character and include those dedicated to photos of “up skirts” (photos taken up a person’s legs in public); contests (e.g., which user-uploaded photos get the most likes); trades (treating victim-survivors like baseball cards); or solicitations (in which users ask for photos of someone); and many others (Citron, 2022; McGlynn et al., 2021; Eaton et al., 2017).
@@ 5. Data Sources
To develop our vignette collection, we performed a literature review focusing on qualitative research related to non-consensual intimate imagery (NCIM). Using the search terms “image-based sexual abuse”, “online sexual abuse”, “sexting”, “revenge pornography”, “intimate partner violence”, and “non-consensual intimate imagery”, we searched ACM Digital Library, ProQuest Research Library, PsycINFO, and Communication & Mass Media Complete databases. The first author selected studies rich in qualitative data that detail the impact of NCIM, prioritizing narratives that provide deep and personal insights into harms. This includes interview studies that documented direct quotes from victim-survivors, and reports that contain stories and quotes from victim-survivors. We excluded the majority of legal reviews and quantitative studies due to the lack of focus on personal narratives. The selected articles were reanalyzed to highlight how specific technologies facilitated and created these narratives and harms. The result is a focused yet rich sample of victim-survivor experiences that illustrate how technology directly enabled NCIM. The sources for the vignettes and quotes are:
@@ 6.2. Applications
The application layer of the sociotechnical stack includes video hosting sites, social media platforms, mobile applications, and online forums. It consists of the rules, policies, and the code that enable these platforms (Lessig, 2009). This layer can facilitate NCIM harms through mechanisms such as impersonation and by providing a space for knowingly hosting non-consensual content.
@@ 6.5.2. Price of storage
Perpetrators of NCIM exploit cheap storage for blackmail and abuse (Bates, 2017; McGlynn et al., 2021). The cost of storing digital content, such as videos and images, is minimal. This affordability means that once someone obtains a copy of content, they are likely to retain it indefinitely. There are little incentives for deletion. Unlike digital storage, physical storage incurs costs. Storing physical photographs in a photo album will ultimately run out of space.
@@ 7. Roadmap for NCIM Research
We use the sociotechnical stack as a framework to identify and explore potential research areas to reduce the harms associated with NCIM. By systematically examining each layer of the sociotechnical stack, we propose targeted interventions that could diminish the technical capabilities that currently facilitate NCIM. This includes innovating on privacy safeguards, enhancing content moderation frameworks, and developing more robust consent mechanisms within digital platforms. This research agenda aims to bridge the gap between technical potential and social necessity, focusing on transforming how social computing and related fields can respond more effectively to challenges of NCIM. No technology will eliminate the harms of NCIM; rather, the intent here is to significantly reduce harms using sociotechnical approaches.
@@ 7.3. Research in algorithms
The emergence of sexually explicit AI-generated content presents another dimension of NCIM harm. Studies indicate that deepfake sexual abuse constitutes a significant portion, accounting for 96 to 98% of all deepfake content (Kwok and Koh, 2021; Karasavva and Noorbhai, 2021; Compton and Hunt, 2024). This alarming prevalence, coupled with existing inadequacies in legal protections, underscores the urgent need for policy interventions. Deepfakes pose unique challenges that redefine the boundaries of consent in the creation of media. Unlike traditional forms of NCIM, deepfakes eliminate the need for physical proximity or explicit recordings, as perpetrators need only easily accessible benign images or videos from the victim’s social media profiles. Exploiting default publicness and leveraging advancements in generative AI to produce nude images, deepfakes facilitate the creation and dissemination of NCIM without consent or awareness. The distinction between consent and violation in deepfake technology is stark. Unlike cases where intimate media was initially shared consensually with a partner, deepfakes leave little room for ambiguity. While not inherently limited to non-consensual content, the most prevalent uses of deepfake technology are harmful, and necessitates a critical reassessment of its legal status and the protections afforded to individuals.
@@ 7.4. Research in networks
Could protocols be utilized to enforce consensual data transmission? Could we then affirm that the transmission of data—particularly intimate content—is consensual? This approach could significantly reduce NCIM harms which primarily stem from unauthorized sharing over networks. However, implementing such a system requires careful governance to balance the deterrence of non-consensual sharing with the need to maintain the free flow of knowledge and other beneficial data. Developing media protocols including licensing agreements could help deter from unauthorized use. One challenge here is applying these solutions to emerging decentralized platforms, where control and enforcement mechanisms are not as straightforward as in centralized systems. Research in this area may consider innovative approaches to protocol design that can operate effectively within different topologies.
@@ 8. Conclusion
NCIM is both a technical and sociotechnical challenge, interacting deeply with many social and computing elements. The roots of sexual abuse and the misogyny that drive NCIM are historical and pre-date the internet. The sociotechnical approach presented in this paper cannot fully solve the underlying issues of abuse, nor fully address the shaming and gendered imbalances experienced by victim-survivors, it offers strategies to reduce the widespread proliferation of harmful content. This paper addresses the critical gap in how technologies not only facilitate but also generate NCIM at scale. Through the sociotechnical stack framework, we analyze the interplay between social impacts and technical elements, highlighting the role of features such as generative AI models in creating non-consensual content. Our approach demonstrated the practical application of the sociotechnical stack to dissect the complexities of NCIM, revealing how these technologies influence the creation, obtainment, and dissemination content. We present our findings and outline a research roadmap of sociotechnical research in NCIM to support victim-survivors, deter perpetrators, and foster consent in digital interactions.

### v3

@@ Abstract.
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; November 09–13, 2024; San José, Costa Ricaprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Collaborative and social computing theory, concepts and paradigms
@@ 1. Introduction
Non-consensual intimate media (NCIM) is the unauthorized creation, sharing, or distribution of sexual content containing someone’s body or likeness. This form of online abuse usually targets women, who constitute about 90% of all cases (Eaton et al., 2017; Initiative, 2014; saf, [n. d.]). NCIM cases can be varied, ranging from sexually explicit deepfakes, to covert recordings, to sextortion (See Fig. 2). NCIM is common—studies suggest 1 in 6 adults have had their intimate content shared without consent (Powell et al., 2024; Ruvalcaba and Eaton, 2020). NCIM is also deeply traumatic—almost all victim-survivors experience severe anxiety and depression. About half seriously consider ending their own lives, and in some cases, women and girls commit suicide to escape emotional pain and social shame. (Compton and Hunt, 2024; Initiative, 2014; Eaton et al., 2017). While there is widespread agreement that NCIM is harmful and some agreement that it should be regulated, there is less consensus on how to do so (Citron and Franks, 2014; De Angeli et al., 2023; Citron and Franks, 2020; Franks, 2014). The law is evolving to recognize NCIM, but access to legal recourse is limited and cannot prevent continued re-distribution. Victim-survivors may request content removal directly from platforms, but these efforts are typically laborious, emotionally taxing, and frequently futile due to the diverse jurisdictions governing online platforms (McGlynn et al., 2021; De Angeli et al., 2023).
@@
This paper highlights the essential links between NCIM and its enabling technical components. Drawing on existing surveys and interviews with more than 400 victim-survivors, we identify a critical gap: technologies, including social media platforms, file formats, and content recommendation algorithms, do not merely enable but actively create NCIM at scale. To systemically analyze this, we introduce a conceptual framework called the sociotechnical stack. In software engineering and computer science, a “tech stack” refers to a layered structure of technologies used to build systems, like the TCP/IP protocols in computing networking. We extend this concept to include social impacts, integrating the technical and social dimensions of systems. Our analysis shows how certain layers, such as a content recommendation algorithm, can unintentionally promote NCIM. The sociotechnical stack illustrates the ways our digital environment facilitates risks and harms for NCIM victim-survivors, including the non-consensual creation of sexual content via generative AI and unauthorized redistribution of privately shared intimate media. We propose a roadmap for sociotechnical research to address NCIM and alleviate risks for victim-survivors.
@@
  • The sociotechnical stack conceptual framework: Intended for analyzing social computing issues, the sociotechnical stack framework breaks down complex sociotechnical problems into specific technical components that contribute to social impacts and harms.
@@ 2. Overview of NCIM
NCIM’s prevalence is extensive and likely underreported, compounded by technical challenges in detection and social stigma that deter reporting. One 2017 survey found that 5% of respondents admitted to non-consensually sharing intimate photos (Eaton et al., 2017). This almost certainly undercounts people who consume NCIM online. A 2014 study estimated around 11,000 websites hosted non-consensual content—a figure that has likely grown (Citron, 2014; Initiative, 2014; Eaton et al., 2017). These sites range in character and include those dedicated to photos of “up skirts” (photos taken up a person’s legs in public); contests (e.g., which user-uploaded photos get the most likes); trades (treating victim-survivors like baseball cards); or solicitations (in which users ask for photos of someone); and many others (Citron, 2022; McGlynn et al., 2021; Eaton et al., 2017).
@@ 5. Data Sources
To develop our vignette collection, we performed a literature review focusing on qualitative research related to non-consensual intimate imagery (NCIM). Using the search terms “image-based sexual abuse”, “online sexual abuse”, “sexting”, “revenge pornography”, “intimate partner violence”, and “non-consensual intimate imagery”, we searched ACM Digital Library, ProQuest Research Library, PsycINFO, and Communication & Mass Media Complete databases. The first author selected studies rich in interview data and first-person accounts that detail the impact of NCIM, prioritizing narratives that provide deep and personal insights into harms. This includes interview studies that documented direct quotes from victim-survivors, and reports that contain stories and quotes from victim-survivors. We excluded the majority of legal reviews and quantitative studies due to the lack of focus on personal narratives. We developed hierarchical codes of the interview text in the published articles. The hierarchical codes later became the themes detailed in Section 6. The selected articles were reanalyzed to highlight how specific technologies facilitated and created these narratives and harms. The result is a focused yet rich sample of victim-survivor experiences that illustrate how technology directly enabled NCIM. The sources for the vignettes and quotes are:
@@ 6.2. Applications
The application layer of the sociotechnical stack includes video hosting sites, social media platforms, mobile applications, and online forums. It consists of the rules, policies, and code that enable these platforms (Lessig, 2009). This layer can facilitate NCIM harms through mechanisms such as impersonation and by providing a space for knowingly hosting non-consensual content.
@@ 6.5.2. Price of storage
Perpetrators of NCIM exploit cheap storage for blackmail and abuse (Bates, 2017; McGlynn et al., 2021). The cost of storing digital content, such as videos and images, is minimal. This affordability means that once someone obtains a copy of content, they are likely to retain it indefinitely. There are few incentives for deletion. Unlike digital storage, physical storage incurs costs. Storing physical photographs in a photo album will ultimately run out of space.
@@ 7. Roadmap for NCIM Research
We use the sociotechnical stack as a framework to identify and explore potential research areas to reduce the harms associated with NCIM. By systematically examining each layer of the sociotechnical stack, we propose targeted interventions that could diminish the technical capabilities that currently facilitate NCIM. This includes innovating privacy safeguards, enhancing content moderation frameworks, and developing more robust consent mechanisms within digital platforms. This research agenda aims to bridge the gap between technical potential and social necessity, focusing on transforming how social computing and related fields can respond more effectively to the challenges of NCIM. No technology will eliminate the harms of NCIM; rather, the intent here is to significantly reduce harms using sociotechnical approaches.
@@ 7.3. Research in algorithms
The emergence of sexually explicit AI-generated content presents another dimension of NCIM harm. Studies indicate that deepfake sexual abuse constitutes a significant portion, accounting for 96 to 98% of all deepfake content (Kwok and Koh, 2021; Karasavva and Noorbhai, 2021; Compton and Hunt, 2024). This alarming prevalence, coupled with existing inadequacies in legal protections, underscores the urgent need for policy interventions. Deepfakes pose unique challenges that redefine the boundaries of consent in the creation of media. Unlike traditional forms of NCIM, deepfakes eliminate the need for physical proximity or explicit recordings, as perpetrators need only easily accessible benign images or videos from the victim’s social media profiles. Exploiting default publicness and leveraging advancements in generative AI to produce nude images, deepfakes facilitate the creation and dissemination of NCIM without consent or awareness. The distinction between consent and violation in deepfake technology is stark. Unlike cases where intimate media was initially shared consensually with a partner, deepfakes leave little room for ambiguity. While not inherently limited to non-consensual content, the most prevalent uses of deepfake technology are harmful, and necessitate a critical reassessment of its legal status and the protections afforded to individuals.
@@ 7.4. Research in networks
Could protocols be utilized to enforce consensual data transmission? Could we then affirm that the transmission of data—particularly intimate content—is consensual? This approach could significantly reduce NCIM harms which primarily stem from unauthorized sharing over networks. However, implementing such a system requires careful governance to balance the deterrence of non-consensual sharing with the need to maintain the free flow of knowledge and other beneficial data. Developing media protocols including licensing agreements could help deter unauthorized use. One challenge here is applying these solutions to emerging decentralized platforms, where control and enforcement mechanisms are not as straightforward as in centralized systems. Research in this area may consider innovative approaches to protocol design that can operate effectively within different topologies.
@@ 8. Conclusion
NCIM is both a technical and sociotechnical challenge, interacting deeply with many social and computing elements. The roots of sexual abuse and the misogyny that drive NCIM are historical and pre-date the internet. The sociotechnical approach presented in this paper cannot fully solve the underlying issues of abuse, nor fully address the shaming and gendered imbalances experienced by victim-survivors, it offers strategies to reduce the widespread proliferation of harmful content. This paper addresses the critical gap in how technologies not only facilitate but also generate NCIM at scale. Through the sociotechnical stack framework, we analyze the interplay between social impacts and technical elements, highlighting the role of features such as generative AI models in creating non-consensual content. Our approach demonstrated the practical application of the sociotechnical stack to dissect the complexities of NCIM, revealing how these technologies influence the creation, obtainment, and dissemination of content. We present our findings and outline a research roadmap of sociotechnical research in NCIM to support victim-survivors, deter perpetrators, and foster consent in digital interactions.
Acknowledgements.
We thank our colleague Daniel Szalkiewicz for early feedback on this work. This material is based upon work supported by the National Science Foundation under Grants 1763297 and 2311102.